# Exchange control in a MOS double quantum dot made using a 300 mm wafer process

### v1

@@ Abstract
Figure 3: |||| Exchange control in the natSi double quantum device. a, Energy diagram depicting two-electron spin states in the (1,1) detuning regime, with pulse sequence steps overlaid. b, Detuning pulse sequence including initialisation to the $\ket{\downarrow\uparrow}$ state via a semi-adiabatic ramp (orange), followed by a non-adiabatic pulse (J) to near zero-detuning to increase the exchange coupling for duration $\tau_{E}$. c, Exchange driven oscillations between $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states, measured rf-phase response proportional to singlet probability, with d, the corresponding Fourier transform. e Ratio between exchange coupling strength $J(\varepsilon)$ and dot-to-dot Zeeman energy difference $\Delta E_{z}$.f Dephasing time $T_{2}^{*}$ extracted from the decay of the exchange oscillations (orange dots) and fit with $T_{2}^{*}=\sqrt{\left<\left(\partial(\hbar\Omega\right)^{2}\right>}/2\hbar$ (purple line) [21, 33]. g Gate quality factor.
@@
$\Delta E_{z}(\phi_{B})/h=\left|B\right|\left|\Delta\alpha-\Delta\beta\sin(2\phi_{B})\right|+g\mu_{B}\sigma_{\text{HF}}.$ (2)
The Rashba and Dresselhaus SOI terms are respectively captured by $\Delta\alpha$ and $\Delta\beta$. We find $\Delta\alpha=5.9\begin{subarray}{c}+1.6\\ -1.5\end{subarray}$z/T and $\Delta\beta=45\begin{subarray}{c}+5\\ -7\end{subarray}~{}$z/T, which are larger than other reported values [18, 40, 45] and could be partially influenced by the large asymmetry in the gate biasing conditions. The fit assumes that for the fixed detuning of $\varepsilon_{P}=0.926$meV used here, the residual exchange interaction $J(\varepsilon_{P})/h=6.3\pm 1.9$MHz is independent of in-plane magnetic field orientation. We operate at an in-plane magnetic field direction near the $\left[1\bar{1}0\right]$ direction at $\phi_{B}=55^{\circ}~{}(235^{\circ})$. Overall, this section expands the recent studies of the SOI in isotopically purified 28Si MOS nanostructures [40, 18, 44, 45] to natural silicon, where the non-negligible effect of the Overhauser field needs to be taken into account.
@@
The extracted $J/\Delta E_{z}$ ratio is shown in Fig. 3, reducing as a function of increasing $\varepsilon$ to a minimum value of $0.5\pm 0.3$ at $\varepsilon=1.054$ meV (beyond this point we cease to observe oscillations). This non-zero minimum shows there remains a residual exchange that cannot be fully turned off, which should be taken into account when designing two-qubit exchange gates.
@@ Acknowledgements
We acknowledge helpful conversations with H. Jnane, A. Seigel, S.C. Benjamin, J.M. Williams, A.J. Fisher and G. Burkard at Quantum Motion. We also acknowledge technical support from G. Antilen Jacob at the London Centre for Nanotechnology. This work received support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 951852 (Quantum Large Scale Integration in Silicon); from the Engineering and Physical Sciences Research Council (EPSRC) under grant Nos. (EP/S021582/1), (EP/L015978/1), (EP/T001062/1) and (EP/L015242/1); and from Innovate UK under grant Nos. (43942) and (10015036). T.M. acknowledges support from the Winton Programme for the Physics of Sustainability. M.F.G.Z. acknowledges support from the UKRI Future Leaders Fellowship (MR/V023284/1).
@@
J.F.C.W. conducted the experiments and analysed the results presented in this work with input from R.C.C.L., M.A.F., M.F.G.Z. and J.J.L.M.. J.F.C.W., M.A.F. and R.C.C.L. conducted preliminary experiments. T.M. and G.A.O developed the quantum dot rf simulator in appendix D. T. M. performed the simulations under the supervision of D. F. W. and M. F. G. Z.. S.M.P. designed the device under the supervision of M.A.F., M.F.G.Z. and J.J.L.M.. J.J. and S.K. fabricated the device under the supervision of B.G.. F.E.v.H. characterised the superconducting resonator under the supervision of M.F.G.Z.. N.J. maintained the experimental setup. J.F.C.W., M.F.G.Z. and J.J.L.M. wrote the manuscript with input from R.C.C.L., M.A.F., N.J. and S.M.P.. M.F.G.Z. and J.J.L.M. conceived and oversaw the experiment.

### v2

@@ Abstract
Figure 3: |||| Exchange control in the natSi double quantum device. a, Energy diagram depicting two-electron spin states in the (1,1) detuning regime, with pulse sequence steps overlaid. b, Detuning pulse sequence including initialisation to the $\ket{\downarrow\uparrow}$ state via a semi-adiabatic ramp (orange), followed by a non-adiabatic pulse (J) to near zero-detuning to increase the exchange coupling for duration $\tau_{E}$. c, Exchange driven oscillations between $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states, measured rf-phase response proportional to singlet probability, with d, the corresponding Fourier transform. e Ratio between exchange coupling strength $J(\varepsilon)$ and dot-to-dot Zeeman energy difference $\Delta E_{z}$.f Dephasing time $T_{2}^{*}$ extracted from the decay of the exchange oscillations (orange dots) and fit with $T_{2}^{*}=\sqrt{2}\hbar/\delta{\left(h\Omega\right)}$ (purple line) [21, 33]. g Gate quality factor.
@@
$\Delta E_{z}/h=\sqrt{{\left(\left|B\right|\left|\Delta\alpha-\Delta\beta\sin(2\phi_{B})\right|\right)}^{2}+{\left(g\mu_{B}\sigma_{\text{HF}}/h\right)}^{2}}.$ (2)
The Rashba and Dresselhaus SOI terms are respectively captured by $\Delta\alpha$ and $\Delta\beta$. We find $\Delta\alpha=6.2\begin{subarray}{c}+1.6\\ -1.5\end{subarray}$z/T and $\Delta\beta=45\begin{subarray}{c}+5\\ -7\end{subarray}~{}$z/T, which are larger than other reported values [18, 40, 45] and could be partially influenced by the large asymmetry in the gate biasing conditions. The fit assumes that for the fixed detuning of $\varepsilon_{P}=0.926$meV used here, the residual exchange interaction $J(\varepsilon_{P})/h=6.3\pm 1.9$MHz is independent of in-plane magnetic field orientation. We operate at an in-plane magnetic field direction near the $\left[1\bar{1}0\right]$ direction at $\phi_{B}=55^{\circ}~{}(235^{\circ})$. Overall, this section expands the recent studies of the SOI in isotopically purified 28Si MOS nanostructures [40, 18, 44, 45] to natural silicon, where the non-negligible effect of the Overhauser field needs to be taken into account.
@@
The extracted $J/\Delta E_{z}$ ratio is shown in Fig. 3(e), reducing as a function of increasing $\varepsilon$ to a minimum value of $0.5\pm 0.3$ at $\varepsilon=1.054$ meV (beyond this point we cease to observe oscillations). This non-zero minimum shows there remains a residual exchange that cannot be fully turned off, which should be taken into account when designing two-qubit exchange gates.
@@ Acknowledgements
We acknowledge helpful conversations with J.M. Williams, H. Jnane, A. Seigel, S.C. Benjamin, A.J. Fisher and G. Burkard at Quantum Motion. We also acknowledge technical support from G. Antilen Jacob at the London Centre for Nanotechnology. This work received support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 951852 (Quantum Large Scale Integration in Silicon); from the Engineering and Physical Sciences Research Council (EPSRC) under grant Nos. (EP/S021582/1), (EP/L015978/1), (EP/T001062/1) and (EP/L015242/1); and from Innovate UK under grant Nos. (43942) and (10015036). T.M. acknowledges support from the Winton Programme for the Physics of Sustainability. M.F.G.Z. acknowledges support from the UKRI Future Leaders Fellowship (MR/V023284/1).
@@
J.F.C.W. conducted the experiments and analysed the results presented in this work with input from R.C.C.L., M.A.F., M.F.G.Z. and J.J.L.M.; J.F.C.W., M.A.F. and R.C.C.L. conducted preliminary experiments; T.M. and G.A.O developed the quantum dot rf simulator in appendix D; T. M. performed and wrote about the simulations under the supervision of D. F. W. and M. F. G. Z.; S.M.P. designed the device under the supervision of M.A.F., M.F.G.Z. and J.J.L.M.; J.J. and S.K. fabricated the device under the supervision of B.G.; F.E.v.H. characterised the superconducting resonator under the supervision of M.F.G.Z.; N.J. maintained the experimental setup; J.F.C.W., M.F.G.Z. and J.J.L.M. wrote the manuscript with input from R.C.C.L., M.A.F., N.J. and S.M.P.; M.F.G.Z. and J.J.L.M. conceived and oversaw the experiment.